# The Tension of Space as Dark Energy: A No Geometric Sequestering Theorem for Minimal Sequesters

### v2

@@ Abstract
We model space as an elastic membrane and identify its uniform tension $T_{s}$ with the vacuum energy density of space. The central result is a no geometric sequestering theorem. In any minimal matter vacuum sequester framework where a global constraint cancels constant contributions from the matter sector by adjusting a global variable, a purely geometric unit operator remains intact. Concretely, a volume term of the form $S_{s}=-T_{s}\int\sqrt{-g}\,d^{4}x$ does not couple to the global variable that enforces the matter cancellation, and it survives unchanged in the local field equations. We establish this result within a covariant effective field theory using the background field method. The analysis shows that the coefficient $T_{s}$ is renormalized by graviton loops and in general by mixed matter–graviton loops and therefore runs with the renormalization scale $\mu$. The running defines an effective cosmological term $T_{s}(\mu)$ that is present even after the matter vacuum energy has been neutralized by the sequester constraint. In particular, we find that $T_{s}(\mu)$ sources $G_{\mu\nu}=-T_{s}(\mu)\,g_{\mu\nu}+\cdots$. The tension term $T_{s}(\mu)$ can then be understood as residual dark energy in the minimal sequestering framework.
@@ 1 Introduction
Late in the 1990s the situation changed. Two supernova projects, the Supernova Cosmology Project [7] and the High-Z Supernova Search Team [6], measured Type Ia supernovae as standardizable candles [^1] Type Ia supernovae were used as standard candles due to their remarkably uniform light curves. and found them to be dimmer than expected in a decelerating model. The inferred expansion is accelerating. Independent evidence from the Cosmic Microwave Background anisotropies by DASI [8], WMAP [9], and Planck [10], and from Baryon Acoustic Oscillations [11, 12], supports this conclusion. For a classic review see [13].
@@
The vacuum contribution that drives late time acceleration is often called dark energy. It is dark because it does not radiate or absorb in a way we can directly detect, and its evidence comes from dynamics such as the acceleration rate. The measured mean energy density today is about $5.249\times 10^{-10}\,\mathrm{J\,m^{-3}}$. A naive quantum field theory estimate of the zero point energy gives a value larger by about $122$ orders of magnitude. This huge discrepancy is the cosmological constant problem. The question it raises is simple to state but hard to answer. What sets the tiny value of the vacuum energy that gravitates on cosmic scales, if the quantum contributions of known fields appear so large in simple estimates?
@@
Many ideas have been explored. Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [54], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [57] and multiple top forms [58]. For broad reviews see [30, 31].
Our goal here is to give a simple and intuitive origin for the cosmological term, and hence for dark energy, using only basic ideas that are compatible with but not restricted to string theory. The key input is the physics of extended objects with tension. We use this as a guide to show that a stress which is proportional to the metric arises naturally when a three dimensional medium with a constant volume tension fills space. We then argue that ordinary space itself can be modeled as such a medium. The resulting stress tensor has the same form as vacuum energy.
We begin with two parallel statements. For an infinitely extended fundamental three brane one finds,
$T_{\mu\nu(3)}=-T_{3}\,g_{\mu\nu}$ (1.10)
where $T_{3}$ is the brane tension with dimensions of energy density. If instead we treat space itself as an elastic three brane with an intrinsic volume tension $T_{s}$, then,
$T_{\mu\nu(s)}=-T_{s}\,g_{\mu\nu}$ (1.11)
Both expressions have the same structure as the vacuum stress tensor discussed earlier. The second statement is stronger conceptually because it identifies space itself as the source of the vacuum energy. We will analyze how a geometric tension of space and a four form sequestering structure interact with the vacuum sector and we will show that while matter vacuum energy can be neutralized by global constraints, a purely geometric NG-like tension cannot be removed. The residual cosmological term is then a geometric constant that we denote by $T_{s}$ and we identify it with Dark Energy.
@@ 2.1 Review of Nambu–Goto Dynamics for Strings and Branes
A fundamental string in $1+1$ dimensions has worldsheet action,
@@
For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [32, 33, 34],
@@
2.2 Stress Tensor of an Infinitely Extended Three Brane
@@
as claimed in Sec. 1. This matches the expected units because $T_{3}$ is an energy density. Comparing with equation (1.5) in the Introduction we see that a uniform three brane tension $T_{3}$ behaves gravitationally like a vacuum energy density $\rho_{\mathrm{vac}}c^{2}$.
2.3 Space as the Three Brane
@@
We now look at the stress tensor that follows from the elastic action of space equation (2.20). On varying $S_{s}$, we obtain,
@@
Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}c^{2}$.
Therefore, the Nambu–Goto structure ensures that a uniform volume tension produces a stress proportional to the metric. For an infinitely extended three brane this gives $T_{\mu\nu(3)}=T_{3}g_{\mu\nu}$. When we elevate space itself to the role of the three brane with intrinsic volume tension $T_{s}$, the same computation yields $T_{\mu\nu(s)}=-T_{s}g_{\mu\nu}$. Hence the cosmological term can be interpreted as the elastic volume tension of space. In the next section we will couple this picture to a four form sequestering framework and show that while matter vacuum energy can be neutralized by global constraints, a geometric volume tension cannot be removed. The residual dark energy is then precisely this $T_{s}$.

### v3

# The Tension of Space as Dark Energy: Dynamics and Phenomenology

@@ Abstract
We propose a phenomenological framework in which the observed late-time dark-energy sector is interpreted as the intrinsic tension of space itself. Motivated by recent observational hints that the dark-energy equation of state may exhibit mild low-redshift evolution, we begin from an intrinsic membrane description of spacetime and show that a uniform space tension contributes to the gravitational field equations with precisely the tensor structure of vacuum energy. We then consider its Dirac–Born–Infeld completion, which naturally introduces a hidden Abelian $U(1)$ gauge sector on the membrane of space. Within this setting, we study a late-time hidden symmetry-breaking transition $U(1)_{h}\to\mathbb{Z}_{n}$, which reorganizes hidden magnetic energy into a coarse-grained flux-tube reservoir. If this reservoir exchanges energy with the dynamical part of the space-tension sector, the effective dark-energy density becomes time dependent and acquires a nontrivial equation of state. At the background level, the model yields a simple proof-of-concept realization of mild running dark energy and admits a transient crossing of the phantom divide. We compare the resulting evolution with a compressed observational benchmark in the effective Chevallier–Polarski–Linder plane. Although the minimal single-reservoir realization does not exactly reproduce the benchmark target, it generates a phenomenologically relevant low-redshift evolution and points toward a late hidden-sector transition together with a substantial hidden defect population. The construction should therefore be viewed not as a complete theory of dark energy, but as a physically transparent proof of concept.
@@ 1 Introduction
Late in the 1990s the situation changed. Two supernova projects, the Supernova Cosmology Project [7] and the High-Z Supernova Search Team [6], measured Type Ia supernovae as standardizable candles [^1] Type Ia supernovae were used as standard candles due to their remarkably uniform light curves. and found them to be dimmer than expected in a decelerating model. The conclusion was that the expansion of space is accelerating. Independent evidence from the Cosmic Microwave Background anisotropies by DASI [8], WMAP [9], and Planck [10], and from Baryon Acoustic Oscillations [11, 12], supports this conclusion. For a classic review see [13].
@@
This vacuum contribution that drives late time acceleration is often called dark energy. It is dark because it does not radiate or absorb in a way we can directly detect, and its evidence comes from dynamics such as the acceleration rate. The measured mean energy density today is about $5.249\times 10^{-10}\,\mathrm{J\,m^{-3}}$. A naive quantum field theory estimate of the zero point energy gives a value larger by about $122$ orders of magnitude. This huge discrepancy is the cosmological constant problem. The question it raises is simple to state but hard to answer. What sets the tiny value of the vacuum energy that gravitates on cosmic scales, if the quantum contributions of known fields appear so large in simple estimates?
@@
Many ideas have been explored. Representative proposals include quintessence [15, 16], k-essence [17, 18], modified gravity [19, 20, 21, 22, 23], the Dvali Gabadadze Porrati braneworld [24, 25], Chaplygin gas [26], holographic dark energy models [27, 28, 29], quantum cosmology [30, 31, 32], dynamical relaxation of the Cosmological Constant via membrane nucleation events of top forms [33] and multiple top forms [34]. For broad reviews see [35, 36].
[…]
If this observational trend persists, then the late-time cosmic acceleration may not be sourced by a perfectly rigid cosmological constant, but rather by an effective vacuum sector that evolves slowly over cosmological time. This motivates the search for physically transparent phenomenological descriptions of dynamical dark energy. In the present work we explore one such possibility. Instead of introducing an ad hoc dark-energy fluid from the outset, we interpret the observed vacuum sector as the intrinsic tension of space itself.
The purpose of the present work is more modest than solving the full cosmological constant problem. We do not attempt to explain from first principles why the absolute vacuum scale is so small. We assume that such a suppression may arise through mechanisms such as vacuum-energy sequestering, Hawking-type Euclidean arguments for the cosmological term, four-form flux discretuum scenarios in which many flux sectors allow an exceptionally small effective $\Lambda$ to emerge, or treating the vacuum as a self-sustained superfluid-like medium that can dynamically adjust its effective vacuum energy toward equilibrium, see for example [30, 31, 32, 42, 43, 44, 45, 46, 47, 48, 49, 51, 50, 52, 53, 54]. Instead, we investigate whether the late-time dark-energy sector that survives after all ultraviolet questions are set aside may admit a physically transparent effective description in terms of a running tension of space. This viewpoint allows one to connect the vacuum-like contribution in Einstein’s equations to an intrinsic geometric property of spacetime, and then to ask how additional hidden-sector structure might perturb that contribution away from exact constancy which may result in a dynamical equation of state for dark energy.
The roadmap of the paper is as follows. In Section 2 we formulate the intrinsic membrane picture in its minimal form. Starting from the standard Nambu–Goto action for extended objects, we specialize to the case in which the three-brane is identified with physical space itself, and show explicitly that a uniform space tension contributes to the field equations with exactly the tensor structure of vacuum energy. We then generalize this intrinsic description to its Dirac–Born–Infeld completion, which naturally introduces a hidden Abelian gauge sector on the membrane of space. In Section 3 we investigate a late-time effective origin for the running of the space tension. We show how a hidden $U(1)_{h}\to\mathbb{Z}_{n}$ breaking can generate a coarse-grained flux-tube reservoir, how the total tension sector may be split into a rigid background piece and a dynamical contribution, and how energy exchange between the two leads to a time-dependent effective dark-energy equation of state. In Section 4 we compare the resulting running-tension evolution with the effective CPL benchmark suggested by current late-time cosmological data, using a compressed Gaussian description of the $(w_{0},w_{a})$ posterior. Finally, in Section 5 we discuss the physical interpretation of the construction, its limitations, and possible directions for future work. A complementary companion work, Ref. [55], develops an alternative late-time realization of the same general picture of dynamical dark energy, in which the residual tension of space is supplemented not by a hidden defect reservoir but by a viscous longitudinal phonon fluid living on the elastic brane.
We now begin with the minimal intrinsic description in which space is treated as an elastic membrane with a Nambu–Goto type volume action.
@@ 2.1 Review of Nambu–Goto Dynamics for Strings and Branes
A fundamental string in $1+1$ dimensions has worldsheet action [56],
@@
For a $d$ brane in (d+1)(d{+}1) dimensions the natural generalization is [57, 58],
@@
2.2 Stress Tensor of an Infinitely Extended $D_{3}$ Brane
@@
Comparing with equation (1.5) in the Introduction we see that a uniform three brane tension $T_{3}$ behaves gravitationally like a vacuum energy density $\rho_{\mathrm{vac}}c^{2}$.
2.3 Space as a $D_{3}$ Brane
@@
We now look at the stress tensor that follows from the elastic action of space, that is equation (2.20). On varying $S_{s}$, we obtain,
@@
Thus the elastic model of space reproduces the vacuum stress tensor with the identification $T_{s}=\rho_{\mathrm{vac}}\,c^{2}$. Equivalently, we have,
$T_{s}=\frac{\Lambda c^{4}}{8\pi G}\,.$ (2.29)
In the present work, we take the next step and identify this intrinsic tension of space directly with the observed late-time vacuum sector. More precisely, $T_{s}$ is not interpreted as the naive bare zero-point energy of all ultraviolet degrees of freedom. Rather, it is understood as the small residual vacuum energy density that remains after the ultraviolet physics responsible for suppressing the enormous microscopic contributions has already done its job as mentioned earlier. Our standpoint here is deliberately agnostic about which ultraviolet mechanism is ultimately realized in nature. We simply take the observed small cosmological term as an infrared input and identify it with the effective volume tension of space itself. In this sense, $T_{s}$ should be viewed as the residual vacuum sector that survives after the ultraviolet vacuum-energy problem has been dealt with, rather than as a first-principles derivation of that suppression within the present model.
Therefore, for $\Lambda\approx 1.09\times 10^{-52}\,\mathrm{m}^{-2}$ [59], equation (2.29) gives the tension of space as,
$T_{s}\approx 5.249\times 10^{-10}\;\mathrm{J/m^{3}}$ (2.30)